If the net investment function is given by I (t) = 100e^ {0.1t} calculate

(a) The capital formation from the end of the second year to the end of the fifth year;

(b) The number of years required before the capital stock exceeds $100 000


1
Expert's answer
2022-03-30T04:46:45-0400

(a) we have to calculate:

"I(5) - I(2) = 100 e^{0.1*5} - 100 e^{0.1*2} = 100(1.649 - 1.221)=42.8"

(b)

"I(t) = 100000\\
100 e^{0.1*t} = 100 000\\
e^{0.1*t} = 1000 | ln(.)\\
0.1*t = 6.908\\
t = 69.08"